Find the midpoint of the line segment joining the points P1 and P2.
P1 = (-1, 6); P2 = (-5, 3)

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Find the midpoint of the line segment joining the points P1 and P2.
P1 = (-1, 6); P2 = (-5, 3)
--------
It's the average of x & y separately.
----
x: (-1-5)/2 = -3
y: (6 + 3)/2 = 9/2
---
MP is (-3,9/2)

Answer by stanbon(75887)

Find the midpoint of the line segment joining the points P1 and P2.
P1 = (-1, 6); P2 = (-5, 3)
---------------
Let the midpoint be (x,y)
x = (-1-5)/2 = -6/2 = -3
y = (6+3)/2 = 9/2 = 4.5
---------------
Ans:: (-3,4.5)
------------
Cheers,
Stan H.
-----------